# Modeling strategies for speech enhancement in the latent space of a neural audio codec

### v1

@@ Abstract
Neural audio codecs (NACs) provide compact latent speech representations in the form of sequences of continuous vectors or discrete tokens. In this work, we investigate how these two types of speech representations compare when used as training targets for supervised speech enhancement. We consider both autoregressive and non-autoregressive speech enhancement models based on the Conformer architecture, as well as a simple baseline where the NAC encoder is simply fine-tuned for speech enhancement. Our experiments reveal three key findings: predicting continuous latent representations consistently outperforms discrete token prediction; autoregressive models achieve higher quality but at the expense of intelligibility and efficiency, making non-autoregressive models more attractive in practice; and encoder fine-tuning yields the strongest enhancement metrics overall, though at the cost of degraded codec reconstruction. The code and audio samples are available online.111sofienekammoun.github.io/SE-NAC-25/
@@ 1 Introduction
Classic SE algorithms worked predominantly in the short-time Fourier transform (STFT) domain, exploiting structure in the time–frequency plane and statistical assumptions about speech and noise [1, 2, 3]. Early deep learning approaches to SE followed the established time–frequency pipeline by estimating time-frequency masks or magnitudes using discriminative training of neural networks [4, 5]. Later, time-domain architectures such as Conv-TasNet [6] showed that learned encoder–separator–decoder pipelines with 1D temporal convolutions can outperform oracle time-frequency masking in speech separation, which removes the need for explicit phase estimation by using trainable filterbanks and time-domain loss functions. The effectiveness of this time-domain approach was also confirmed for SE [7].
@@ 2.2.1 Discrete token representation
We first consider modeling strategies where the noisy and clean speech signals are represented as tokens in the discrete latent domain: $\mathbf{x}=\{x_{t,n}\}_{t=1,n=1}^{T,N}$, $\mathbf{y}=\{y_{t,n}\}_{t=1,n=1}^{T,N}\in\{1,...,K\}^{N\times T}$.
@@ 2.2.2 Continuous latent representation
We now consider modeling strategies where the noisy and clean speech signals are represented as continuous latent vectors, i.e., $\bar{\mathbf{x}}=\{\bar{\mathbf{x}}_{t}\}_{t=1}^{T},\bar{\mathbf{y}}=\{\bar{\mathbf{y}}_{t}\}_{t=1}^{T}\in\mathbb{R}^{L\times T}$. So, at a time step $t$, instead of predicting a categorical distribution over the discrete tokens for each quantization stage $n$, the model will directly provide a prediction of the continuous vector $\bar{\mathbf{x}}_{t}$, given $\bar{\mathbf{y}}$ and potentially additional conditions. Additionally, the input embedding layer becomes a simple linear layer that projects input vectors from dimension $L$ to $H$.
@@ 2.3 Training and inference
All models considered here are trained in a supervised fashion on paired noisy–clean speech data by maximizing the conditional likelihood of clean speech given noisy speech, which is simply defined by taking the logarithm of equations (2) to (5). When the model is defined over the continuous representation $\bar{\mathbf{x}}$, maximizing the likelihood is equivalent to minimizing the mean squared error (MSE). In the other case, when the model is defined over the discrete latent representation $\mathbf{x}$, this amounts to minimizing the cross-entropy loss.
Independent of the representation space, given a trained model, SE is achieved by taking the argmax over $\mathbf{x}$ or $\bar{\mathbf{x}}$ of the above-defined probabilistic models. In this work, we are not interested in introducing stochasticity in the SE process by sampling the distributions. We would rather estimate the most likely clean speech signal for a given trained model and noisy signal. For continuous-prediction models, the estimated vectors are quantized before being passed to the NAC decoder, ensuring compatibility with the codec reconstruction pipeline, and for discrete-prediction models, the predictions go through the detokenization process described in (1).
@@ 3 Experiments
Model OVRL ($\uparrow$) SIG ($\uparrow$) BAK ($\uparrow$) P808 ($\uparrow$) UTMOS ($\uparrow$) CosSim ($\uparrow$) dWER ($\downarrow$) GFLOPs ($\downarrow$)
DCCRNet [28] 2.80 3.15 4.03 3.54 3.01 96.6 11.80 26
DCUNet [29] 2.97 3.33 3.96 3.62 3.07 96.7 10.11 250
ConvTasNet [6] 3.11 3.39 4.01 3.31 3.27 96.4 11.30 10
DPTNet [30] 3.00 3.32 4.05 3.31 3.38 96.6 10.75 2
AnCoGen [31] 3.00 3.32 4.05 3.31 3.38 96.6 19.30 -
D-AR 2.90 3.17 3.99 3.53 2.76 95.2 25.09 5857
D-NAR 2.89 3.18 3.94 3.50 2.72 95.5 23.12 6
@@ 3.1 Experimental setup
Dataset   Libri1Mix [32] is a single-speaker noisy speech dataset constructed from clean utterances of LibriSpeech [33] combined with noise samples from WHAM! [34]. It covers a range of signal-to-noise ratio (SNR) values between 6-6 and $3$ dB, simulating challenging noisy conditions. We use the train-360 subset for our experiments, which provides $156$ hours of paired noisy/clean speech for training and $4$ hours each for validation and testing.
@@ 3.2 Results
To further investigate the weakness of discrete models, we introduce a variant of D-NAR that uses continuous inputs $\bar{\mathbf{y}}$ instead of discrete noisy tokens $\mathbf{y}$. This allows us to measure to what extent the input representation affects the D-NAR model. This model (denoted by D-NAR in Table 1) performs better than other discrete variants, but still lags behind continuous models, suggesting that conditioning on discrete tokens is only part of the problem and that the main bottleneck lies in the output space and associated loss function used for discrete prediction.
Additionally, whether continuous or discrete, AR models tend to achieve higher quality (DNSMOS, UTMOS) than NAR models, likely due to their ability to capture temporal dependencies during generation. However, they demonstrate degraded speaker similarity and intelligibility, which could be explained by error accumulation over autoregressive steps. Moreover, the quality gains of continuous AR over continuous NAR are modest compared to the significant computational overhead and the higher dWER values.
Finally, we experiment with fine-tuning the NAC encoder alongside training a C-NAR enhancement model. We chose the C-NAR system to explore this fine-tuning strategy because it already provides a strong balance between quality and efficiency. This strategy, denoted by C-NAR-FT in Table 1, yields the best compromise between speech quality, intelligibility, and inference speed.

### v2

@@ Abstract
Neural audio codecs (NACs) provide compact latent speech representations in the form of sequences of continuous vectors or discrete tokens. In this work, we investigate how these two types of speech representations compare when used as training targets for supervised speech enhancement. We consider both autoregressive and non-autoregressive speech enhancement models based on the Conformer architecture, as well as a simple baseline where the NAC encoder is simply fine-tuned for speech enhancement. Our experiments reveal three key findings: predicting continuous latent representations consistently outperforms discrete token prediction; autoregressive models achieve higher quality but at the expense of intelligibility and efficiency, making non-autoregressive models more attractive in practice; and adding encoder fine-tuning yields the strongest enhancement metrics overall, though at the cost of degraded codec reconstruction. The code and audio samples are available online.111sofienekammoun.github.io/SE-NAC-25/
@@ 1 Introduction
Classic SE algorithms worked predominantly in the short-time Fourier transform (STFT) domain, exploiting structure in the time–frequency plane and statistical assumptions about speech and noise [1]. Early deep learning approaches to SE followed the established time–frequency pipeline by estimating time-frequency masks or magnitudes using discriminative training of neural networks [21, 32]. Later, time-domain architectures such as Conv-TasNet [19] showed that learned encoder–separator–decoder pipelines with 1D temporal convolutions can outperform oracle time-frequency masking in speech separation, which removes the need for explicit phase estimation by using trainable filterbanks and time-domain loss functions. The effectiveness of this time-domain approach was also confirmed for SE [6].
@@ 2.2.1 Discrete token representation
We first consider modeling strategies where the noisy and clean speech signals are represented as discrete tokens extracted using a pretrained NAC: $\mathbf{x}=\{x_{t,n}\}_{t=1,n=1}^{T,N}\in\{1,...,K\}^{N\times T}$, $\mathbf{y}=\{y_{t,n}\}_{t=1,n=1}^{T,N}\in\{1,...,K\}^{N\times T}$.
@@ 2.2.2 Continuous latent representation
We now consider modeling strategies where the noisy and clean speech signals are represented as continuous latent vectors extracted from a pretrained NAC, i.e., $\bar{\mathbf{x}}=\{\bar{\mathbf{x}}_{t}\}_{t=1}^{T},\bar{\mathbf{y}}=\{\bar{\mathbf{y}}_{t}\}_{t=1}^{T}\in\mathbb{R}^{L\times T}$. So, at a time step $t$, instead of predicting a categorical distribution over the discrete tokens for each quantization stage $n$, the model will directly provide a prediction of the continuous vector $\bar{\mathbf{x}}_{t}$, given $\bar{\mathbf{y}}$ and potentially additional conditions. Additionally, the input embedding layer becomes a simple linear layer that projects input vectors from dimension $L$ to $H$.
@@ 2.3 Training and inference
All models considered here are trained in a supervised fashion on paired noisy–clean speech data by maximizing the conditional likelihood of clean speech given noisy speech, which is simply defined by taking the logarithm of equations (2) to (5). When the model is defined over the continuous representation $\bar{\mathbf{x}}$, maximizing the likelihood is equivalent to minimizing the mean squared error (MSE). In the other case, when the model is defined over the discrete latent representation $\mathbf{x}$, this amounts to minimizing the cross-entropy loss. For both AR models, teacher forcing is used during training. For the C-AR model, the input vectors $\bar{\mathbf{x}}_{1:t-1}$ are quantized using the NAC’s RVQ module before feeding them to the Conformer model, in order to mitigate the problem of error accumulation that usually occurs for AR models.
Independent of the representation space, given a trained model, SE is achieved by taking the argmax over $\mathbf{x}$ or $\bar{\mathbf{x}}$ of the above-defined probabilistic models. In this work, we are not interested in introducing stochasticity in the SE process by sampling the distributions. We would rather estimate the most likely clean speech signal for a given trained model and noisy signal. For continuous-prediction models, the estimated vectors are quantized before being passed to the NAC decoder, ensuring compatibility with the codec reconstruction pipeline, and for discrete-prediction models, the predictions go through the detokenization process described in (1). For simplicity, we segment noisy speech audio into chunks of $1$ second and process them separately for all of our experiments.
@@ 3 Experiments
Model OVRL ($\uparrow$) SIG ($\uparrow$) BAK ($\uparrow$) P808 ($\uparrow$) UTMOS ($\uparrow$) CosSim ($\uparrow$) dWER ($\downarrow$) GFLOPs ($\downarrow$) Parameters (M)
DCCRNet [10] 2.80 3.15 4.03 3.54 3.01 96.6 11.80 26 3.7
DCUNet [3] 2.97 3.33 3.96 3.62 3.07 96.7 10.11 250 7.7
ConvTasNet [19] 3.11 3.39 4.01 3.31 3.27 96.4 11.30 10 5.0
DPTNet [2] 3.00 3.32 4.05 3.31 3.38 96.6 10.75 2 2.8
AnCoGen [26] 3.00 3.32 4.05 3.31 3.38 96.6 19.30 - -
D-AR 2.90 3.17 3.99 3.53 2.76 95.2 25.09 5857 82.2
D-NAR 2.89 3.18 3.94 3.50 2.72 95.5 23.12 6 68.7
@@ 3.1 Experimental setup
Dataset   Libri1Mix [5] is a single-speaker noisy speech dataset constructed from clean utterances of LibriSpeech [22] combined with noise samples from WHAM! [33]. It covers a range of signal-to-noise ratios (SNRs) from 6-6 to $3$ dB, simulating challenging noisy conditions. We use the train-360 subset for training, which contains $156$ hours of paired noisy/clean speech, and the test and dev subsets for validation and testing, which each contain $4$ hours.
@@ 3.2 Results
To further investigate the weakness of discrete models, we introduce a variant of D-NAR that uses continuous inputs $\bar{\mathbf{y}}$ instead of discrete noisy tokens $\mathbf{y}$. This allows us to measure to what extent the input representation affects the D-NAR model. This model (denoted by D-NAR in Table 1) performs better than other discrete variants, but still lags behind continuous models, suggesting that conditioning on discrete tokens is only part of the problem and that the main bottleneck lies in the output space and associated loss function used for discrete prediction. Additionally, whether continuous or discrete, AR models tend to achieve higher quality (DNSMOS, UTMOS) than NAR models, likely due to their ability to capture temporal dependencies during generation. However, they demonstrate degraded speaker similarity and intelligibility, which could be explained by error accumulation over autoregressive steps. Moreover, the quality gains of continuous AR over continuous NAR are modest compared to the significant computational overhead and the higher dWER values.
We also experiment with fine-tuning the NAC encoder alongside training a C-NAR enhancement model. We chose the C-NAR system to explore this fine-tuning strategy because it already provides a strong balance between quality and efficiency. This strategy, denoted by C-NAR-FT in Table 1, yields the best compromise between speech quality, intelligibility, and inference speed.
@@
Finally, we perform a last experiment in which the C-NAR model is trained to predict a time-frequency mask from the STFT of the noisy speech signal. All other model hyperparameters remain the same as before, including the architecture and the MSE loss function. As reported in Table 1, this STFT-NAR model yields the weakest performance across all metrics (except dWER), confirming the advantage of working in the NAC’s latent space.